OpenStudy (anonymous):

-2+7i ------ -4-3i A.) 3-14i --------- 6 B.) -13-34i --------- 25 C.) -13-22i --------- 25 D.) -13+22i ---------- 25

OpenStudy (amistre64):

xavier is quick like lol

OpenStudy (anonymous):

Multiply the top and bottom by -4+3i. I usually only give advice when im all ninja like

OpenStudy (amistre64):

-16-9=-25... got that much lol

OpenStudy (amistre64):

16-9 tho right? 7

OpenStudy (amistre64):

ack.... 25 is the bottom

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

okay top is -27 and bottom is11

OpenStudy (amistre64):

16 -9i^2 = 16+9 = 25

OpenStudy (anonymous):

How did you get that katie?

OpenStudy (anonymous):

ohh wait its 25 sorry

OpenStudy (amistre64):

27 is a magic number lol

OpenStudy (anonymous):

so is the answer D

OpenStudy (anonymous):

I got B: (-2+7i)(-4+3i)=8-6i-28i+21i^2=8-6i-28i-21=-13-34i For the numerator. 25 Denominator

OpenStudy (anonymous):

oh okay i see how you did that

OpenStudy (anonymous):

i gave you a medal

OpenStudy (anonymous):

for your help

OpenStudy (amistre64):

I get B as well